OpenStudy (anonymous):

Samantha gets an electricity bill.The cost of the electricity used before the VAT is added is 49.30 pounds. VAT is charged at 5% on domestic fuel bills.What is the cost of the electricity bill,including VAT?

OpenStudy (anonymous):

493.80 pounds-correction

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (michele_laino):

the 5% of VAT is computed on the amount 49.30, so the requested cost is: \[\Large 49.30 + \left( {49.30 \cdot \frac{5}{{100}}} \right) = ...?\]

OpenStudy (anonymous):

sorry 493.80

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} 49.30 + \left( {49.30 \cdot \frac{5}{{100}}} \right) = ...? \hfill \\ \hfill \\ = 49.3 + 2.465 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

with the correct of 493.80 following what you have written my answer came out as 24.69

OpenStudy (michele_laino):

please wait, I update my computation...

OpenStudy (anonymous):

so do I just add the value of 24.69 to 493.80?

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

\[\Large \begin{gathered} 493.80 + \left( {493.80 \cdot \frac{5}{{100}}} \right) = ...? \hfill \\ \hfill \\ = 493.80 + 24.69 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

518.49

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thanks! can you please help me with the next question i will tag you.

OpenStudy (michele_laino):

ok!